mhchen:

Linear Algebra: Matricies with negative exponents: Let A be a matrix. Does A^(-x) = [A^(-1)]^(x) ?

mhchen:

\[A = \left[\begin{matrix}a & b \\ c & d\end{matrix}\right], A^{-2}=(A^{-1})^{2} \] In other words, is this true

mhchen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mhchen \[A = \left[\begin{matrix}a & b \\ c & d\end{matrix}\right], A^{-2}=(A^{-1})^{2} = (A^{-1})*(A^{-1})\] In other words, is this true \(\color{#0cbb34}{\text{End of Quote}}\)